OpenStudy (anonymous):

Help, help??(: Given triangle DEF with D(-4,-1), E(-1,8), and F(5,4), find the median DT in point-slope form.

OpenStudy (ranga):

The median joins the vertex D to the mid point of the opposite side EF. Mid point of EF = ((-1+5)/2, (8+4)/2) The median passes trough the midpoint and (-4,-1) You know two point on the line and you can compute the equation of the line using the formula: (y-y1) = m(x-x1) where m = (y2-y1)/(x2-x1)

OpenStudy (anonymous):

Wait I don't get that point-slope formula part? What do I plug in where?

OpenStudy (anonymous):

find the coordinates of point t

OpenStudy (anonymous):

But what's point T??

OpenStudy (anonymous):

@Peter14

OpenStudy (anonymous):

the median or something? I don't exactly know.

OpenStudy (anonymous):

It's confusing .-. hah thanks anyways

OpenStudy (anonymous):

 that would be my guess, but it might not be right

OpenStudy (anonymous):

That's what I thought too..

OpenStudy (anonymous):

so T is exactly between E and F. find the coordinates, then find the slope of the line DT as ranga showed, and use either D or T as the point for point-slope form

OpenStudy (anonymous):

oooooh, ok thanks (:

OpenStudy (ranga):



OpenStudy (ranga):

T = ((-1+5)/2, (8+4)/2) T = (2,6) Slope of line DT is: difference in y-coordinates / difference in x-coordinates D = (-4,-1) ; T = (2,6) slope m = (6-(-1)) / (2-(-4)) = 7/6 Point-Slope form is: (y - y1) = m(x - x1) The line passes through points D and T. Let us pick the point D (-4,-1). Plug m = 7/6, x1 = -4, y1 = -1 (y - (-1)) = 7/6(x - (-4)) y + 1 = 7/6(x + 4) Had we picked point T (2,6) the equation would be: y - 6 = 7/6(x - 2) The equations may look different but if you multiply it out they are the same. If you have multiple choice answers pick the one they provide. The equation in the slope-intercept form is: y = 7/6x + 11/3 (but they don't ask for this).

OpenStudy (anonymous):

Ooh, thank youu!!!!